OpenStudy (anonymous):

For the squence an = 3n-5 what is the value of a30

OpenStudy (anonymous):

a1 = 3.1 - 5 = 3 - 5 = -2 a2 = 3.2 - 5 = 6 - 5 = 1 a3 = 3.3 - 5 = 9 - 5 = 4 a4 = 3.4 - 5 = 12 - 5 = 7 ... find a30?

OpenStudy (anonymous):

You can do either by calculating utp 30 or see this sequence -2, 1, 4,7, .... is this reminds you of Arithemetic progression.